OpenStudy (anonymous):

help me solve 14-4y < (with the line) 66 a) y >(with the line ) -13 b) y >-17 c) y> (with the line)-17 d)y >-13

ganeshie8 (ganeshie8):

add 4y to both sides

OpenStudy (johnweldon1993):

\[14 - 4y \le66\] add 4y to both sides.....then subtract 66 from both sides....and finally divide both sides by 4...what do you get?

OpenStudy (johnweldon1993):

We do this to isolate 'y' because that is what we are solving for \[14 - 4y \le 66\] after adding 4y to both sides we have \[14 \le 66 + 4y\] then we subtract 66 from both sides \[-52 \le 4y\] and we finally divide both sides by 4 to completely isolate 'y' \[\frac{ -52 }{ 4 } \le \frac{ 4y }{ 4 }\] simplified down a bit \[y \ge \frac{ -52 }{ 4 }\] so what does that equate to?

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

so its a

OpenStudy (johnweldon1993):

Right it simplifies down to \[y \ge -13\] So yes...answer choice A :)

OpenStudy (anonymous):

thank you (:

OpenStudy (johnweldon1993):

anytime! :)